OpenStudy (twizttiez):

Complete the equation for the linear function whose graph contains the point (−1, 2) and has a slope of 0.9. y = x +

OpenStudy (danjs):

same thing.. use the point-slope formula y - y1 = m*(x-x1) point(x1,y1) slope m

OpenStudy (danjs):

just put the values into the equation.

OpenStudy (twizttiez):

y = 0.9x + 2.9 ah hah!

OpenStudy (danjs):

y - 2 = 0.9* ( x - (-1) )

OpenStudy (danjs):

y - 2 = 0.9x + 0.9 y = 0.9x + 0.9 + 2

OpenStudy (twizttiez):

So i was wrong again D: no!!!

OpenStudy (danjs):

using the point slope form, you get y - 2 = 0.9 *(x + 1) then distribute the right, and add 2 to both sides y - 2 = 0.9x + 0.9 y = 0.9x + 2.9

OpenStudy (twizttiez):

Gah Dan i was right you fooled me curses!

OpenStudy (danjs):

oh, i didn't see that, i was looking for the precess to get to that.. sorry

OpenStudy (twizttiez):

No prob xD